OpenStudy (anonymous):

A triangle with an area of 100 cm² has a height of 10 cm. What is the triangle's base movement? a.20 cm b. 1000 cm c.10 cm d.5 cm

jimthompson5910 (jim_thompson5910):

A = (bh)/2 100 = (b*10)/2 100 = (10b)/2 100*2 = 10b 200 = 10b 200/10 = b 20 = b b = 20 So the base is 20 cm

OpenStudy (anonymous):

the answer is a :) 20cm

OpenStudy (anonymous):

Thanks both:)